### v2

# IR/UV mixing from higher-order interactions in a Scalar Field

Abstract
The mismatch between the observed vacuum energy and the value suggested by quantum-field-theoretic estimates is a foundational puzzle. While Lorentz-invariant regularization softens the naive quartic divergence, it still leaves a discrepancy of many orders of magnitude and appears incompatible with Weinberg’s no-go theorem, which prohibits obtaining a small cosmological constant within local, translationally invariant quantum field theory (QFT) without fine tuning. One broad class of ideas that can evade these constraints involves nonlocal or quasi-nonlocal assumptions that induce UV/IR limit mixing.
In this work I present a concrete four-dimensional scalar field theory with a quasi-local, marginal (by power-counting) four-field interaction whose momentum-dependent term generates a dominant $k^{8}$ contribution to the effective Hamiltonian of each Fourier mode at large wave number $k$. When the field is confined to a finite spacetime box of side $L$, the resulting dynamics render each high-$k$ mode an effective quartic oscillator rather than a harmonic one. A semiclassical estimate of the ground-state energy then enforces an emergent UV cutoff
$k_{\rm cutoff}\;\propto\;k_{\rm Pl}^{3/8}\,k_{\rm box}^{5/8}$
where $k_{\rm Pl}$ is a microscopic UV scale, usually chosen to be the Planck scale while $k_{\rm box}\sim 2\pi/L$ is set by the size of the region. For cosmological-scale boxes this cutoff is dramatically reduced relative to $k_{\rm Pl}$, yielding a strong suppression of the naive vacuum-energy density, while for laboratory-sized boxes $k_{\rm cutoff}\approx k_{\rm Pl}$, leaving standard QFT phenomenology intact.
I discuss stability, the approximations used, and generalizations to other higher-order interactions. I then compute the resulting vacuum-energy density using dimensional regularization and propose a possible way to deduce the scaling of the effective equation of state with $L$. I then include qualitative comparison with the DESI indication of mildly time-varying dark energy. The aim is to provide a field-theoretic mechanism illustrating how quasi-local interactions can generate UV/IR mixing and dynamically modify the UV cutoff scale.
I Introduction
The discrepancy between naive quantum-field-theoretic estimates of the vacuum energy and the observed value inferred from cosmology spans dozens of orders of magnitude. The most familiar form of this discrepancy arises from summing zero-point energies of free fields up to a UV cutoff $\Lambda$:
$\rho_{\rm vac}^{\rm naive}\sim\frac{1}{2}\int^{\Lambda}\!\frac{d^{3}\vec{k}}{(2\pi)^{3}}\,\omega_{k}\;\sim\;\Lambda^{4}$ (1)
with $\omega_{k}=({\vec{k}}^{2}+m^{2})^{1/2}$. Choosing $\Lambda$ near the Planck scale overshoots the observed vacuum energy by about 120 orders of magnitude. Lorentz-invariant regularization schemes such as dimensional regularization preserve $p=-\rho$ for the vacuum state and significantly reduce the result, yet still leave a discrepancy Akhmedov ; Koksma ; Jerome . Weinberg’s no-go theorem Weinberg further shows that achieving a small cosmological constant in local, translationally invariant QFT typically requires unwanted fine tuning.
These issues motivate exploring mechanisms in which the UV behavior is dependent on the behavior in the IR. UV/IR mixing is familiar in non-commutative field theories Minwalla , effective theories with gravitational consistency bounds CKN1 ; CKN2 , and various nonlocal modifications of gravity Anupam . The key theme is that even small nonlocal proposals can profoundly reorganize the relationship between UV and IR scales.
In this work I propose a simple, explicit scalar-field model that exhibits a UV/IR-mixed (effective) cutoff when analyzed in a finite box. The approach is conservative and is meant to show a conceptual framework:
  • the interaction is quasi-local but chosen for calculational transparency rather than derivation from a fundamental theory;
  • approximations are stated clearly, including where manifest Lorentz invariance is lost;
  • the suggested cosmological implications are exploratory.
The central observation is that a marginal (by power-counting) four-field interaction with a momentum-dependent kernel yields a dominant $k^{8}$ term in each mode’s Hamiltonian at high $k$, converting the harmonic-oscillator Hamiltonian into a quartic oscillator. A semiclassical estimate of the ground-state energy in this regime then enforces an emergent cutoff dependent on both the Planck scale and the IR box size.
The structure of the paper is as follows. Section II reviews the standard vacuum-energy estimate. Section III defines the quasi-local interaction in position space and Section IV rewrites it in momentum space. Section V explains the controlled approximations used. Section VI derives the effective Hamiltonian and its quartic form. Section VII treats stability and Ostrogradsky issues. Section VIII derives the emergent UV/IR-mixed cutoff. Section IX analyzes generalizations. Section X analyzes how this theory might apply to fermions. Section XI computes the vacuum-energy density using dimensional regularization. Section XII discusses cosmological implications. Conclusions appear in Section XIII.
II The Naive Calculation
As explained in Jerome , the naive calculation just computes the energy of the vacuum states. This is done for a standard free scalar field whose Lagrangian density is given by
${\cal L}=\bigg[\frac{1}{2}\partial^{\mu}\phi(x)\>\partial_{\mu}\phi(x)-\frac{1}{2}m^{2}\phi(x)^{2}\bigg]$ (2)
[…]
To set the stage for what follows, we will define two Fourier transforms (remember $k$ is a discrete set of variables, as the field is confined to a box of space-time dimension $L$).
$\hat{\phi}(k)=\frac{1}{V_{4}}\sum_{k}\>\phi(x)\>e^{ik.x}\>\>\>\>\>,\>\>\>\>\>\phi^{*}(k)=\phi(-k)$ (3)
where $k\equiv(\vec{k},\omega)$ is the 4-momentum vector and $k^{2}=\omega^{2}-\vec{k}^{2}$. And, in addition
$\tilde{\phi}(\vec{k},t)=\frac{1}{V_{3}}\sum_{\vec{k}}\>\phi(\vec{x},t)\>e^{i\vec{k}.\vec{x}}\>\>\>\>\>,\>\>\>\>\>\phi^{*}(\vec{k},t)=\phi(-\vec{k},t)$ (4)
By inspection, the mass dimension of $\hat{\phi}(k)$ is 3-3, while that of $\tilde{\phi}(\vec{k},t)$ is 2-2.
With these definitions, we can now write the action integral as
${\cal S}=\int d^{4}x\>{\cal L}=\frac{1}{V_{4}}\sum_{k}\>\frac{(k^{2}-m^{2}){\hat{\phi}(k)}{\hat{\phi}(-k)}}{2}\>\>\>\>\>\>\>\>\>\>\>$ (5)
$\equiv\int dt\>L=\frac{1}{V_{3}}\int dt\>\sum_{\vec{k}}\bigg(\frac{1}{2}{\dot{\tilde{\phi}}}(\vec{k},t){\dot{\tilde{\phi}}}(-\vec{k},t)\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>$
$-\frac{\vec{k}^{2}+m^{2}}{2}{\tilde{\phi}}(\vec{k},t){\tilde{\phi}}(-\vec{k},t)\bigg)\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>$
The Hamiltonian for the field is written from the final term for the Lagrangian as
$H=\frac{1}{V_{3}}\>\sum_{\vec{k}}\bigg(\frac{1}{2}{\dot{\tilde{\phi}}}(\vec{k},t){\dot{\tilde{\phi}}}(-\vec{k},t)+\frac{\vec{k}^{2}+m^{2}}{2}{\tilde{\phi}}(\vec{k},t){\tilde{\phi}}(-\vec{k},t)\bigg)$ (6)
We can now quantize the field, using the standard prescriptions, where we write
$\phi(\vec{x},t)=\sum_{\vec{k}}\frac{1}{\sqrt{2\omega_{k}}}\>\bigg(a(\vec{k})e^{i\vec{k}.\vec{x}-i\omega_{k}t}+a^{\dagger}(\vec{k})e^{i\vec{k}.\vec{x}+i\omega_{k}t}\bigg)$ (7)
and we obtain the usual result
$H=\sum_{\vec{k}}\omega_{k}\>\bigg(a^{\dagger}(\vec{k})a(\vec{k})+\frac{1}{2}\bigg)$ (8)
where $\omega_{k}=\sqrt{\vec{k}^{2}+m^{2}}$.
The naive argument for the vacuum energy per unit volume in the model consists of simply summing up the energies of every mode in the ground state, leading to the integral
${\cal E}=\sum_{\vec{k}}\frac{1}{2}\omega_{k}\approx\int_{|k|=0}^{|k|=\Lambda}\frac{d^{3}\vec{k}}{(2\pi)^{3}}\sqrt{\vec{k}^{2}+m^{2}}\>\sim\>\Lambda^{4}$ (9)
where the regularization cut-off for the integral over wave-vectors is usually chosen at the wave-vector corresponding to a mode whose energy is the Planck energy, i.e., wave-vector $k_{Pl}\approx\sim 10^{35}m^{-1}$ corresponding to $E_{Pl}\approx 10^{19}GeV$. The result crucially depends on the computation that the energy of the mode is roughly proportional to the wave-vector and hence the cut-off wave-vector is as large as it is, i.e., $k_{Pl}$.
III Quasi-local four-field interaction
We now introduce a power-counting–marginal four-field interaction built from derivative operators and a rapidly decaying kernel. In the below, the $\Box$ operator is the usual d’Alembertian:
$S_{\rm int}$ $=-g\!\!\prod_{i=1}^{4}\!\!\int\!d^{4}x_{i}\;(\Box\phi)(x_{1})(\Box\phi)(x_{2})(\Box\phi)(x_{3})(\Box\phi)(x_{4})$
$\hskip 2.84544pt\times\,\delta^{(4)}(x_{4}+x_{1}-x_{2}-x_{3})\,C(x_{1}-x_{2})C(x_{1}-x_{3})$
$\hskip 5.69046pt\times C(x_{2}-x_{3})C(2x_{1}-x_{2}-x_{3})$ (10)
The function $C(x)$ is chosen to be dimensionless, entire, and rapidly decaying; for definiteness we take
$C(x)=\exp[-A^{4}(x.x)^{2}]$ (11)
though alternatives with similar decay (such as $C(x)=\exp[-A^{4}(x.x)^{2}]\Theta(x^{2})$) yield the same qualitative result and preserve causality. The scale $A$ sets the range of the quasi-locality.
The overall interaction is marginal by power counting if $C(x)$ has mass dimension zero. The coupling $g$ is taken to be $O(1)$.
IV Momentum space resolution of the added term
Expanding the scalar field in Fourier modes,
$\phi(x)=\frac{1}{V_{4}}\sum_{k}\tilde{\phi}(k)e^{ik\cdot x}\>\>,\>\>\qquad k=(\omega,\vec{k})$ (12)
with spacetime volume $V_{4}=L^{4}$, and denoting by $\widehat{C}(q)$ the Fourier transform of $C(x)$, one can rewrite Eq. (10) in momentum space. The rapidly decaying, entire nature of $C(x)$ implies that $\widehat{C}(q)$ is smooth and approaches a constant for small $q$. For the present purpose, using this small-momentum approximation,
$\widehat{C}(q)\approx\widehat{C}(0)\equiv\frac{Q}{A^{4}}$ (13)
is adopted, where $Q$ is a positive numerical constant.
After performing the integrals over the positions $x_{i}$ and the Fourier variables associated with the kernels, the interaction simplifies to
$S_{\rm int}=-g\,\frac{Q^{4}}{(AL)^{16}}\sum_{k}k^{8}\,\tilde{\phi}(k)\tilde{\phi}(-k)\tilde{\phi}(-k)\tilde{\phi}(k)$ (14)
It may be useful to note that the real-space term is a convolution that was constructed as such to produce this result. Up to normalization conventions, this may be viewed as a quartic interaction with a $k^{8}$-dependent coefficient for each Fourier mode.
V Summary of the Approximations
Three approximations are used and are listed below:
  1. 1.
    We approximate $\widehat{C}(q)\approx\widehat{C}(0)=Q/A^{4}$ for small momenta $q$; this is justified because $C(x)$ is rapidly decaying and entire.
  2. 2.
    We replace higher time derivatives by their leading contributions in the high-$k$ regime, where spatial derivatives dominate. We use this to derive an approximate (and not manifestly Lorentz invariant) Hamiltonian.
  3. 3.
    After these steps the effective Lagrangian is not manifestly Lorentz invariant; this is acceptable because: (i) the underlying interaction is Lorentz invariant, and (ii) the effective truncation is used only to estimate the ground-state energy scaling relevant for cutoff determination.
VI The full Lagrangian, Hamiltonian and Quartic-Oscillator Form
The total Lagrangian we obtain
${\cal S}=\frac{1}{V_{4}}\bigg(\sum_{k}\>\frac{(k^{2}-m^{2}){\hat{\phi}(k)}{\hat{\phi}(-k)}}{2}-g\frac{{\cal Q}^{4}}{(AL)^{16}}\sum_{k}\>(k^{2})^{4}{\hat{\phi}}(k){\hat{\phi}}(-k){\hat{\phi}}(-k){\hat{\phi}}(k)\bigg)$ (15)
We note that this is Lorentz invariant and the vacuum state it generates is also Lorentz invariant Koksma . This observation will be useful in the section on the computation of the vacuum energy.
Additionally, combining the free Hamiltonian and the interaction term (and dropping time derivatives, as mentioned in point (2) in the previous section) yields
$H=\frac{1}{V_{3}}\sum_{\bm{\vec{k}}}\left[\frac{1}{2}|\dot{\tilde{\phi}}(\bm{\vec{k}},t)|^{2}+\frac{1}{2}(\vec{k}^{2}+m^{2})|\tilde{\phi}(\bm{\vec{k}},t)|^{2}+R_{k}|\tilde{\phi}(\bm{\vec{k}},t)|^{4}\right]$ (16)
with
$R_{k}=\frac{gQ^{4}L^{2}{\vec{k}}^{8}}{(LA)^{16}}$ (17)
For large $k$, the quartic term dominates, giving an effective quartic oscillator.
VII Analysis of potential instabilities
To study a simplified situation, we look at the non-local Lagrangian per mode $\bf L$}, keeping dimensions accurate. For each mode, let’s set $x(t)=\frac{{\tilde{\phi}}(\vec{k},t)}{L^{2}}$, so that $x(t)$ is the dimensionless coordinate corresponding to each mode. Then, we consider Equation (15) in this notation as
${\cal S}=\int dt\sum_{\vec{k}}L\bigg(\frac{1}{2}x^{*}(t)(-\partial_{t}^{2}-\frac{1}{2}\omega_{\vec{k}}^{2})x(t)-\frac{g{\cal Q}^{4}}{4(AL)^{16}}L^{4}f(t)\int dt_{2}\int dt_{2}f^{*}(t_{2})f^{*}(t_{3})f(t_{2}+t_{3}-t)\bigg)$
$\rightarrow{\bf L}=\frac{L}{2}{\dot{x}}(t){\dot{x}}^{*}(t)-\frac{L\Omega^{2}}{2}x(t)x^{*}(t)\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>$
$-\frac{rL^{5}}{4}f(t)\int dt_{2}\int dt_{3}f^{*}(t_{2})f^{*}(t_{3})f(-t+t_{2}+t_{3})\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>$ (18)
where $f(t)=(-{\ddot{x}}-ax)$, $a={\vec{k}}^{2}$ and the above toy Lagrangian is applicable to each mode separately. We will study this Lagrangian; note that the time integrals are over the time dimension $L$ and $x$ is dimensionless, while $\bf L$} has dimensions of $M^{1}$. For convenience, we have used the notation $r=\frac{g{\cal Q}^{4}}{(AL)^{16}}$ and assume that $g\sim 1,{\cal Q}\sim 1,(AL)\sim 1$, so that $r\approx 1$. We write down the Euler-Lagrange equation, varying $x^{*}$, which would be considered independent of $x$ and get
$L{\ddot{x}}^{*}+L\Omega^{2}{\dot{x}}^{*}+\frac{rL^{5}}{2}(-\frac{d^{2}}{dt^{2}}-a)\int dt_{2}\int dt_{3}f(t_{2})f^{*}(t_{3})f(t+t_{3}-t_{2})=0$ (19)
The solution of the above equation with $r=0$ is simple harmonic motion with angular frequency $\Omega$. Now, let’s add the $r$ terms back and study a solution of the form $x(t)=\epsilon e^{i\omega t}$. We obtain the equation
$\omega^{2}=\Omega^{2}+\frac{rL^{6}}{2}(\omega^{2}-a)^{4}\epsilon^{2}$ (20)
We seem to have run into a problem here, this equation has complex solutions for $\omega$. Let us simplify the discussion by setting $\epsilon=1$ and define $\kappa=\frac{rL^{6}}{2}$. We will assume $r\sim 1$ as before. Note that the equation, which is essentially
$\omega^{2}=\Omega^{2}+\kappa(\omega^{2}-a)^{4}$ (21)
has the property that $\kappa^{-\frac{1}{6}}\equiv\Omega_{\kappa}$ is an extremely tiny frequency ($\sim L^{-1}$ in mass units and we will apply this computation to a box the size of the universe, i.e., $L\sim\frac{1}{10^{-42}\>GeV}$) compared to any reasonable masses of scalar particles (unless they are actually 0). We could write the equation, using $a=k^{2}$ as
$\omega^{2}=\Omega^{2}+\frac{(\omega^{2}-k^{2})^{4}}{\Omega_{\kappa}^{6}}$ (22)
It is possible to solve this iteratively and we get approximately
$\omega^{2}\approx k^{2}+\Omega_{\kappa}^{3/2}(k^{2}-\Omega^{2})^{1/4}$
$\omega\approx k+\frac{\Omega_{\kappa}^{3/2}}{2k}(k^{2}-\Omega^{2})^{1/4}$ (23)
If the term $(k^{2}-\Omega^{2})$ were to be negative, the imaginary part of $\omega$ would be proportional to $\frac{\Omega_{\kappa}^{3/2}\Omega^{1/2}}{k}\propto\frac{L^{-3/2}\Omega^{1/2}}{k}$ if $\Omega>k$.
If we apply this to a box the size of the Universe, for particles whose masses are of order $\Omega\sim 100\>GeV$ and $k\sim 10^{-19}GeV$ (which we will compute later), we get a tiny imaginary part for $\omega$, roughly $10^{-42}GeV$. This means that one e-folding will take as long as the life of the Universe, hence we choose to neglect this imaginary part.
We conclude that the dispersion relation does not produce physically relevant unstable behavior.
Additionally, given this Lagrangian L, and assuming that $\omega$ represents the energy of the state with wave-vector $k$, the Hamiltonian H can be computed. We find
${\bf L}=L\frac{\omega^{2}-\Omega^{2}}{2}\epsilon^{2}-\frac{r\>L^{7}\epsilon^{4}}{4}(\omega^{2}-a)^{4}$
${\bf H}=L\frac{\omega^{2}+\Omega^{2}}{2}\epsilon^{2}+\frac{r\>L^{7}\epsilon^{4}}{4}(\omega^{2}-a)^{4}$ (24)
The Hamiltonian is clearly positive definite.
Since this is a reasonably oscillatory system, we can address the point made in point (2) in Section $V$, that we could drop the ${\ddot{\tilde{\phi}}}(\vec{k},t)$ term compared to $\vec{k}^{2}{\tilde{\phi}}(\vec{x},t)$. Let us study this point in the quartic regime of this oscillator. That is because ${\ddot{\tilde{\phi}}}\propto\phi^{3}$ for a quartic oscillator and expectation values of even powers of $\phi^{3}$ are smaller than expectation values of even powers of $\phi$ in the ground state of such an oscillator - the excursions from the center are smaller. To say this differently, in the quartic regime, the amplitude of the oscillation is not a free parameter and it is inversely proportional to the square root of the magnitude of the forcing coefficient. Hence, expectation values of higher even powers of the displacement are smaller. Basically
$\ddot{x}=-Ux^{3}\>\>\>\>\>\rightarrow\>\>\>\>\>x\sim A\rightarrow A\sim\sqrt{\frac{1}{U}}$ (25)
Since the forcing factor is proportional to $k^{8}$ in our quartic oscillator, the amplitude is smaller for large $k$. This makes higher even powers of $x$ have smaller expectation values.
VIII The Hamiltonian and the Emergent UV/IR-Mixed Cutoff
The Hamiltonian is as in Equation (16),
$H=\frac{1}{V_{3}}\>\sum_{\vec{k}}\bigg(\frac{1}{2}|{\dot{\tilde{\phi}}}(\vec{k},t)|^{2}+\frac{\omega_{k}^{2}}{2}|{\tilde{\phi}}(\vec{k},t)|^{2}+{\cal R}_{k}|{\tilde{\phi}}(\vec{k},t))|^{4}\bigg)$ (26)
where ${\cal R}_{k}=g\frac{{\cal Q}^{4}}{(LA)^{16}}\>L^{2}\>\>(\vec{k})^{8}$. The Hamiltonian can be checked to have mass dimension $1$. The mass dimension of ${\cal R}_{k}$ can be calculated to be 6.
In the regime where the quartic term dominates for a given mode, the Hamiltonian takes the form of a quartic oscillator. A simple semiclassical estimate Biswas of the ground-state energy for such an oscillator, based on the Bohr–Sommerfeld quantization condition, yields
${\cal E}_{0}(k)\simeq\frac{1}{8^{2/3}}[R_{k}V_{3}]^{1/3}$ (27)
where $V_{3}=L^{3}$ is the spatial volume. Substituting $R_{k}$ gives (we define the box wave-vector $k_{Box}=\frac{2\pi}{L}$)
${\cal E}_{0}(k)\approx\frac{1}{8^{2/3}}({\cal R}_{k}V_{3})^{\frac{1}{3}}=\lambda\bigg(\frac{k^{8}}{k_{Box}^{5}}\bigg)^{\frac{1}{3}}$ (28)
where $\lambda$ is a positive numerical constant of order unity. Requiring ${\cal E}_{0}(k_{\rm cutoff})\sim k_{\rm Pl}$ leads to
$k_{\rm cutoff}\;\sim\;k_{\rm Pl}^{3/8}k_{\rm box}^{5/8}$ (29)
This relation exhibits the desired UV/IR mixing: the effective UV cutoff in wave number is not set solely by the microscopic scale $k_{\rm Pl}$ but also depends on the IR scale $k_{\rm box}$ determined by the size of the spacetime region. The expression can be cast in the schematic form
${\cal E}_{0}(k)\sim k_{\rm Pl}\left(\frac{k}{k_{\rm cutoff}}\right)^{8/3}$ (30)
where $k_{\rm cutpff}$ is the emergent cutoff scale that depends on both $k_{\rm Pl}$ and $k_{\rm box}$.
For a box size comparable to the observable universe, one may take $k_{\rm box}\sim 10^{-23}\,{\rm m}^{-1}$, while $k_{\rm Pl}\sim 10^{35}\,{\rm m}^{-1}$. In this case the formula above gives $k_{\rm cutoff}\sim 10^{-1}\,{\rm m}^{-1}$, corresponding to a very small energy scale. In contrast, for laboratory-scale boxes with $k_{\rm box}$ much larger, $k_{\rm cutoff}$ approaches $k_{\rm Pl}$ and standard QFT phenomenology is preserved.
Figure 1: Emergent cutoff $k_{\mathrm{cutoff}}$ as a function of box length $L$.
Figure 2: Vacuum–energy suppression ratio $\rho_{\mathrm{eff}}/\rho_{\mathrm{naive}}$ vs. $L$.
This corresponds to a linear dispersion energy $10^{-8}$ eV, which is a little smaller than that corresponding to the observed cosmological constant $\sim 10^{4}m^{-1}$. Cutting the usual “naive” integral off at this level leads to a vacuum energy estimate that is $10^{136}$ times smaller than the “naive” estimate and $\sim 10^{60}$ times smaller than the more reasonable estimate Koksma ; Jerome . Note that we get this result purely from (non-linear) terms that respect the original symmetries of the Lagrangian.
We next consider the possibility that higher order terms of this kind might be more dominant. We explore that in the next section.
IX Generalizations
The interaction constructed in Eq. (10) is only one representative of a broader class of quasi-local interactions. Consider more generally interaction terms of the form
$S_{\rm int}^{(q,n)}\sim\frac{g}{M_{\rm Pl}^{s}}\prod_{i=1}^{2n-1}\int d^{4}x_{i}\,(\Box^{q}\phi(x_{i}))\,\Box^{q}\phi(x_{2n})\,\prod_{\alpha}C(P_{\alpha})$ (31)
where $q$ is a non-negative integer, $n$ is the number of field pairs, $M_{\rm Pl}$ is a reference mass scale, the $P_{\alpha}$ denote independent combinations of position differences, and the product over $\alpha$ runs over $Z$ such combinations. Power counting, together with the mass dimension of $C$, fixes $s$.
A representative sample of parameter choices, assuming $A=k_{\rm box}$ for definiteness, is summarized in Table 1. Only the case $q=1$, $n=2$, and dimensionless $C$ produces a low emergent cutoff of the type discussed in the previous section; other choices lead to either very large cutoffs or scales that are not phenomenologically useful for vacuum-energy suppression.
$q$ $n$ $Z$ mass dim. of $C$ $k_{\rm cut}$ (m-1)
1 2 4 0 $10^{-1}$
1 2 4 1 $10^{27}$
1 2 4 2 $10^{56}$
1 2 4 3 $10^{85}$
2 3 15 0 $10^{11}$
2 3 15 1 $10^{47}$
2 3 15 2 $10^{83}$
Table 1: Representative emergent cutoffs for generalized quasi-local interactions, assuming $A=k_{\rm box}$. Only the first line (the model analyzed in detail in this work) leads to a very low cutoff for cosmological boxes.
X Fermions
Fermionic oscillators can have only restricted types of quartic or higher terms and all those terms would be zero in the ground state, as the states would have zero occupation number. More precisely, these terms are of the form $(c^{\dagger}_{i}c_{i})(c^{\dagger}_{j}c_{j})$ where $i,j$ are quantum states. Since, in the vacuum state, these states would be unpopulated, such terms would not contribute.
Hence, the computation for fermions would simply restrict the integration to values of the momentum wave-vector smaller than that computed in the above.
XI Vacuum energy computation
Given the modified dispersion relation encoded in the ground-state energy ${\cal E}_{0}(k)$, the vacuum energy density can be computed using dimensional regularization. In $d$ spatial dimensions, the expression
$\rho_{\rm vac}=\frac{\mu^{4-d}}{2}\int\!\!\frac{d^{d}\bm{k}}{(2\pi)^{d}}\,{\cal E}_{0}(k)$ (32)
is considered, where $\mu$ is an arbitrary mass scale and ${\cal E}_{0}(k)$ is taken in a form that interpolates between the modified large-$k$ behavior and the infrared regime. A convenient choice is
${\cal E}_{0}(k)=k_{\rm Pl}\left[\frac{(k^{2}+k_{\rm cut}^{2})^{4}}{k_{\rm cut}^{8}}\right]^{1/3}$ (33)
Performing the integral in dimensional regularization and continuing to $d=3$ one finds
$\rho_{\rm vac}\;=\;-\frac{0.42}{16\pi^{3/2}}\,k_{\rm Pl}k_{\rm cut}^{3}$ (34)
where the numerical coefficient arises from ratios of Gamma functions.
Substituting $k_{\rm Pl}\sim 10^{19}$ GeV and $k_{\rm cut}\sim 10^{-17}$ GeV yields a vacuum-energy density $\rho_{\rm vac}\sim 10^{-34}\,{\rm GeV}^{4}$, which is substantially reduced compared to the naive estimate and closer to the observed dark-energy scale. We can also repeat the calculations in Akhmedov ; Koksma ; Jerome for all the particle species and obtain $\rho_{vac}^{TOTAL}\approx 3\times 10^{-33}GeV^{4}$, a positive number. Note, again, this $k_{cutoff}$ is very sensitive to the product $(AL)$, as described earlier in the paper.
XII Exploring Cosmological Implications
The discussion so far has been framed in flat spacetime with a finite box of side $L$. In a cosmological setting, the relevant spatial region expands with the scale factor $a(t)$, so that $L\propto a(t)$ and $k_{\rm box}\propto a^{-1}$. The emergent cutoff then scales as
$k_{\rm cutoff}\propto k_{\rm box}^{5/8}\propto a^{-5/8}$ (35)
Since $\rho_{\rm vac}\propto k_{\rm Pl}\,k_{\rm cutoff}^{3}$, one obtains
$\rho_{\rm vac}\propto a^{-15/8}$ (36)
Treating this vacuum contribution as a fluid with energy density $\rho$ and pressure $p$ in a Friedmann–Robertson–Walker background leads to the continuity equation
$\dot{\rho}+3H(\rho+p)=0$ (37)
where $H=\dot{a}/a$ is the Hubble parameter. Writing $p=w\rho$ and using $\rho\propto a^{-15/8}$, one finds
$w=-1-\frac{1}{3}\frac{d\ln\rho}{d\ln a}=-\frac{3}{8}$ (38)
The resulting equation-of-state parameter is not consistent with the nearly $w\simeq-1$ dark energy favored by current observations. However, the mild evolution rate
$\frac{\dot{\rho}}{\rho}=-\frac{15}{8}\frac{\dot{a}}{a}=-\frac{15}{8}H\approx-1.8\,H$ (39)
is of the same order as the deviations from a pure cosmological-constant behavior hinted at by recent DESI results, which suggest a mildly time-varying dark-energy density. A proper cosmological treatment would require evolving the full nonlocal theory in an expanding background, which lies beyond the scope of this exploratory analysis.
XIII Conclusions and further steps
A scalar-field theory with a quasi-local four-field interaction has been constructed that exhibits UV/IR mixing. The interaction generates a $k^{8}$-dependent quartic term in each mode’s Hamiltonian, and for sufficiently large $k$ the mode behaves as a quartic oscillator whose ground-state energy scales with $k$ in a way that enforces an emergent cutoff
$k_{\rm cutoff}\sim k_{\rm Pl}^{3/8}k_{\rm box}^{5/8}$ (40)
This dynamically lowered cutoff suppresses the vacuum-energy density for cosmological-size boxes while leaving ordinary particle-physics scales unaffected in laboratory-sized regions.
The analysis here is intended to be restricted in many ways to illustrate the method, rather than propose a fundamental theory. The kernel $C(x)$ is chosen phenomenologically; approximations are made that break manifest Lorentz invariance; and the cosmological discussion is only qualitative. Nevertheless, the model demonstrates how quasi-local interactions in four-dimensional QFT can produce UV/IR mixing and significantly reduce the effective UV scale that enters vacuum-energy estimates.
An important direction for further work is to study such interactions in fully covariant nonlocal frameworks. It would then be useful to examine their compatibility with precision cosmology and other observational constraints.
XIV Acknowledgments
The hospitality of the Rutgers Physics & Astronomy Department and the NHETC is gratefully acknowledged. Useful discussions with Scott Thomas are also acknowledged.